OpenStudy (anonymous):

In your own words explain the difference between the following two expressions: (-1)^2/4 and (-1)^1/2

OpenStudy (anonymous):

\((-1)^\frac{2}{4}\) \((-1)^\frac{1}{2}\) this is what do u mean ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:) so whats ur opinian ? any diffrence ?

OpenStudy (anonymous):

I don't think so.

OpenStudy (anonymous):

2/2=1 and 4/2=2 which equals -1^1/2

OpenStudy (anonymous):

\((-1)^\frac{2}{4}=\sqrt[4]{(-1)^2}\) \((-1)^\frac{1}{2}=\sqrt[2]{(-1)}\) nw can u see the diffrence ?

OpenStudy (anonymous):

Yes, but how would I explain the difference??

OpenStudy (anonymous):

simply first one =1 secondone = i=sqrt -1 so just say in ur words wat u got